# On the semigroup of monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with the two-element family $\mathscr{F}$ of inductive nonempty subsets of $\omega$

### v1

@@ Abstract.
We study the semigroup of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with the two-elements family $\mathscr{F}$ of inductive nonempty subsets of $\omega$. We describe the structure of elements of the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
Bicyclic monoid, inverse semigroup, bicyclic extension, monoid endomorphism, non-injective, Green’s relations, left-zero semigroup, direct product.
@@
Let $\mathscr{P}(\omega)$ be the family of all subsets of $\omega$. For any $F\in\mathscr{P}(\omega)$ and $n\in\mathbb{Z}$ we put $nF=\{nk\colon k\in F\}$ if $F\neq\varnothing$ and $n\varnothing=\varnothing$. A subfamily $\mathscr{F}\subseteq\mathscr{P}(\omega)$ is called ${\omega}$-closed if $F_{1}\cap(-n+F_{2})\in\mathscr{F}$ for all $n\in\omega$ and $F_{1},F_{2}\in\mathscr{F}$. For any $a\in\omega$ we denote $[a)=\{x\in\omega\colon x\geqslant a\}$.
A subset $A$ of $\omega$ is said to be inductive, if $i\in A$ implies $i+1\in A$. Obvious, that $\varnothing$ is an inductive subset of $\omega$.
@@
    By Lemma 6 from [4] nonempty subset $F\subseteq\omega$ is inductive in $\omega$ if and only $(-1+F)\cap F=F$.
@@
    Since the set $\omega$ with the usual order is well-ordered, for any nonempty inductive subset $F$ in $\omega$ there exists nonnegative integer $n_{F}\in\omega$ such that $[n_{F})=F$.
@@
It is well-known that every automorphism of the bicyclic monoid $\boldsymbol{B}_{\omega}$ is the identity self-map of $\boldsymbol{B}_{\omega}$ [1], and hence the group $\mathbf{Aut}(\boldsymbol{B}_{\omega})$ of automorphisms of $\boldsymbol{B}_{\omega}$ is trivial. In [8] it is proved that the semigroup $\mathrm{\mathbf{End}}(\boldsymbol{B}_{\omega})$ of all endomorphisms of the bicyclic semigroup $\boldsymbol{B}_{\omega}$ is isomorphic to the semidirect products $(\omega,+)\rtimes_{\varphi}(\omega,*)$, where $+$ and *** are the usual addition and the usual multiplication on $\omega$.
@@
This paper is a continuation of [7]. We study non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. We describe the structure of elements of the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
By Proposition 1 of [5] for any $\omega$-closed family $\mathscr{F}$ of inductive subsets in $\mathscr{P}(\omega)$ there exists an $\omega$-closed family $\mathscr{F}^{*}$ of inductive subsets in $\mathscr{P}(\omega)$ such that $*[0)\in\mathscr{F}^{*}$ and the semigroups $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and $\boldsymbol{B}_{\omega}^{\mathscr{F}^{*}}$ are isomorphic. Hence without loss of generality we may assume that the family $\mathscr{F}$ contains the set $[0)$.
@@
Since $\mathfrak{e}$ is an monoid endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$, $(0,0,[0))\mathfrak{e}=(0,0,[0))$. By Proposition 3 of [4] the subsemigroup $\boldsymbol{B}_{\omega}^{\{[0)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the bicyclic semigroup and hence by Corollary 1.32 of [1] the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ either is isomorphic to the bicyclic semigroup or is a cyclic subgroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. If $S$ is a subsemigroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ which is isomorphic to the bicyclic semigroup then by Proposition 4 of [5] there exists $F\in\mathscr{F}$ such that $S\subseteq\boldsymbol{B}_{\omega}^{\{F\}}$. Since $(0,0,[0))\mathfrak{e}=(0,0,[0))$, Proposition 4 from [5] implies that $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$ in the case when the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ is isomorphic to the bicyclic semigroup. In the case when the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ is isomorphic to the cyclic group we have that the equality $(0,0,[0))\mathfrak{e}=(0,0,[0))$ implies that $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\{(0,0,[0))\}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$, because by Theorem 4$(iii)$ of [4] every $\mathscr{H}$-class in $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a singleton.
@@
By $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ we denote the semigroup of all non-injective monoid endomorphisms of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$.
Theorems 1 and 2 describe the algebraic structure of the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
$(k,k,[0))=(1,1,[0))\mathfrak{e}\preccurlyeq(s,s,[0))=(0,0,[1))\mathfrak{e}\preccurlyeq(0,0,[0))=(0,0,[0))\mathfrak{e}$
@@
$(1,0,[1))\mathfrak{e}=((0,1,[1))^{-1})\mathfrak{e}=((0,1,[1))\mathfrak{e})^{-1}=(x,y,[0))^{-1}=(y,x,[0)).$
@@
$(s,s,[0))$ $=(0,0,[1))\mathfrak{e}=((0,1,[1))\cdot(1,0,[1)))\mathfrak{e}=(0,1,[1))\mathfrak{e}\cdot(1,0,[1))\mathfrak{e}=$
$=(x,y,[0))\cdot(y,x,[0))=(x,x,[0))$
@@
$(s+p,s+p,[0))$ $=(1,1,[1))\mathfrak{e}=((1,0,[1))\cdot(0,1,[1)))\mathfrak{e}=(1,0,[1))\mathfrak{e}\cdot(0,1,[1))\mathfrak{e}=$
$=(y,x,[0))\cdot(x,y,[0))=(y,y,[0)).$
@@
$(0,1,[1))\mathfrak{e}=(s,s+p,[0))\quad\hbox{and}\quad(1,0,[1))\mathfrak{e}=(s+p,s,[0)).$
@@
$(0,n_{1},[1))\mathfrak{e}=(\underbrace{(0,1,[1))\cdot\ldots\cdot(0,1,[1))}_{n_{1}\hbox{\tiny{-times}}})\mathfrak{e}=\underbrace{(0,1,[1))\mathfrak{e}\cdot\ldots\cdot(0,1,[1))\mathfrak{e}}_{n_{1}\hbox{\tiny{-times}}}=(s,s+p,[0))^{n_{1}}=(s,s+n_{1}p,[0))$
@@
$(n_{2},0,[1))\mathfrak{e}=(\underbrace{(1,0,[1))\cdot\ldots\cdot(1,0,[1))}_{n_{2}\hbox{\tiny{-times}}})\mathfrak{e}=\underbrace{(1,0,[1))\mathfrak{e}\cdot\ldots\cdot(1,0,[1))\mathfrak{e}}_{n_{2}\hbox{\tiny{-times}}}=(s+p,s,[0))^{n_{2}}=(s+n_{2}p,s,[0)),$
@@
$(k(m+1),k(m+1),[0))=(m+1,m+1,[0))\mathfrak{e}$ $\preccurlyeq(s+pm,s+pm,[0))=(m,m,[1))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(m,m,[0))\mathfrak{e}=(km,km,[0)).$
@@
$(i,j,[0))\gamma_{k_{1}}\gamma_{k_{2}}=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[0))\gamma_{k_{1}}\delta_{k_{2}}=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[0))\delta_{k_{1}}\gamma_{k_{2}}=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[1))\delta_{k_{1}}\gamma_{k_{2}}=(k_{1}(i+1),k_{1}(j+1),[0))\gamma_{k_{2}}=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
@@
$(i,j,[0))\delta_{k_{1}}\delta_{k_{2}}=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[1))\delta_{k_{1}}\delta_{k_{2}}=(k_{1}(i+1),k_{1}(j+1),[0))\delta_{k_{2}}=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
@@
By $\mathfrak{e}_{\boldsymbol{0}}$ we denote the annihilating monoid endomorphism of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$, i.e., $(i,j,[p))\mathfrak{e}_{\boldsymbol{0}}=(0,0,[0))$ for all $i,j\in\omega$ and $p=0,1$. We put $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})=\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})\setminus\{\mathfrak{e}_{\boldsymbol{0}}\}$. Theorem 2 implies that $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ is a subsemigroup of $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
    $\mathfrak{e}_{1}\mathscr{D}\mathfrak{e}_{2}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if there exists a positive integer $k$ such that $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\{\gamma_{k},\delta_{k}\}$;
@@
Since $\mathfrak{e}_{\boldsymbol{0}}$ is zero of the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, the classes of equivalence of Green’s relations of non-zero elements of $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ in the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ coincide with their corresponding classes of equivalence in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, and moreover we have that $\boldsymbol{L}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{R}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{H}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{D}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{J}_{\mathfrak{e}_{\boldsymbol{0}}}=\{\mathfrak{e}_{\boldsymbol{0}}\}$ in the semigroup $\boldsymbol{End}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.

### v2

# On the semigroup of monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with a two-element family $\mathscr{F}$ of inductive nonempty subsets of $\omega$

@@ Abstract.
We study the semigroup of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with a two-elements family $\mathscr{F}$ of inductive nonempty subsets of $\omega$. We describe the structure of elements of the semigroup $\boldsymbol{End}^{*}_{0}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product of the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
Bicyclic monoid, inverse semigroup, bicyclic extension, monoid endomorphism, non-injective, Green’s relations, left-zero semigroup, direct product
@@
Let $\mathscr{P}(\omega)$ be the family of all subsets of $\omega$. For any $F\in\mathscr{P}(\omega)$ and $n\in\mathbb{Z}$ we put $n+F=\{n+k\colon k\in F\}$ if $F\neq\varnothing$ and $n+\varnothing=\varnothing$. A subfamily $\mathscr{F}\subseteq\mathscr{P}(\omega)$ is called ${\omega}$-closed if $F_{1}\cap(-n+F_{2})\in\mathscr{F}$ for all $n\in\omega$ and $F_{1},F_{2}\in\mathscr{F}$. For any $a\in\omega$ we denote $[a)=\{x\in\omega\colon x\geqslant a\}$.
A subset $A$ of $\omega$ is said to be inductive, if $i\in A$ implies $i+1\in A$. Obviously, $\varnothing$ is an inductive subset of $\omega$.
@@
    By Lemma 6 from [4] a nonempty subset $F\subseteq\omega$ is inductive in $\omega$ if and only $(-1+F)\cap F=F$.
@@
    Since the set $\omega$ with the usual order is well-ordered, for any nonempty inductive subset $F$ in $\omega$ there exists a nonnegative integer $n_{F}\in\omega$ such that $[n_{F})=F$.
@@
It is well-known that every automorphism of the bicyclic monoid $\boldsymbol{B}_{\omega}$ is the identity self-map of $\boldsymbol{B}_{\omega}$ [1], and hence the group $\mathbf{Aut}(\boldsymbol{B}_{\omega})$ of automorphisms of $\boldsymbol{B}_{\omega}$ is trivial. In [8] it is proved that the semigroup $\mathrm{\mathbf{End}}(\boldsymbol{B}_{\omega})$ of all endomorphisms of the bicyclic semigroup $\boldsymbol{B}_{\omega}$ is isomorphic to the semidirect products $(\omega,+)\rtimes_{\varphi}(\omega,*)$, where $+$ and * are the usual addition and the usual multiplication on $\omega$.
@@
This paper is a continuation of [7]. We study non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. We describe the structure of elements of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
By Proposition 1 of [5] for any $\omega$-closed family $\mathscr{F}$ of inductive subsets in $\mathscr{P}(\omega)$ there exists an $\omega$-closed family $\mathscr{F}^{*}$ of inductive subsets in $\mathscr{P}(\omega)$ such that $[0)\in\mathscr{F}^{*}$ and the semigroups $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and $\boldsymbol{B}_{\omega}^{\mathscr{F}^{*}}$ are isomorphic. Hence without loss of generality we may assume that the family $\mathscr{F}$ contains the set $[0)$.
@@
By Proposition 3 of [4] the subsemigroup $\boldsymbol{B}_{\omega}^{\{[0)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the bicyclic semigroup and hence by Corollary 1.32 of [1] the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ either is isomorphic to the bicyclic semigroup or is a cyclic subgroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. Since $(0,0,[0))\mathfrak{e}=(0,0,[0))$, Proposition 4 from [5] implies that $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$ in the case when the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ is isomorphic to the bicyclic semigroup. In the other case we have that the equality $(0,0,[0))\mathfrak{e}=(0,0,[0))$ implies that
$(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\{(0,0,[0))\}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}},$
because by Theorem 4$(iii)$ of [4] every $\mathscr{H}$-class in $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a singleton.
@@
By $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ we denote the semigroup of all non-injective monoid endomorphisms of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$.
Theorems 1 and 2 describe the algebraic structure of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
$(k,k,[0))$ $=(1,1,[0))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(s,s,[0))=$
$=(0,0,[1))\mathfrak{e}\preccurlyeq$
[…]
$=(0,0,[0))\mathfrak{e}$
@@
$(1,0,[1))\mathfrak{e}$ $=((0,1,[1))^{-1})\mathfrak{e}=$
$=((0,1,[1))\mathfrak{e})^{-1}=$
$=(x,y,[0))^{-1}=$
$=(y,x,[0)).$
@@
$(s,s,[0))$ $=(0,0,[1))\mathfrak{e}=$
$=((0,1,[1))\cdot(1,0,[1)))\mathfrak{e}=$
$=(0,1,[1))\mathfrak{e}\cdot(1,0,[1))\mathfrak{e}=$
$=(x,y,[0))\cdot(y,x,[0))=$
$=(x,x,[0))$
@@
$(s+p,s+p,[0))$ $=(1,1,[1))\mathfrak{e}=$
$=((1,0,[1))\cdot(0,1,[1)))\mathfrak{e}=$
$=(1,0,[1))\mathfrak{e}\cdot(0,1,[1))\mathfrak{e}=$
$=(y,x,[0))\cdot(x,y,[0))=$
$=(y,y,[0)).$
@@
$(0,1,[1))\mathfrak{e}=(s,s+p,[0))\qquad\hbox{and}\qquad(1,0,[1))\mathfrak{e}=(s+p,s,[0)).$
@@
$(0,n_{1},[1))\mathfrak{e}$ $=(\underbrace{(0,1,[1))\cdot\ldots\cdot(0,1,[1))}_{n_{1}\hbox{\tiny{-times}}})\mathfrak{e}=$
$=\underbrace{(0,1,[1))\mathfrak{e}\cdot\ldots\cdot(0,1,[1))\mathfrak{e}}_{n_{1}\hbox{\tiny{-times}}}=$
$=(s,s+p,[0))^{n_{1}}=$
$=(s,s+n_{1}p,[0))$
@@
$(n_{2},0,[1))\mathfrak{e}$ $=(\underbrace{(1,0,[1))\cdot\ldots\cdot(1,0,[1))}_{n_{2}\hbox{\tiny{-times}}})\mathfrak{e}=$
$=\underbrace{(1,0,[1))\mathfrak{e}\cdot\ldots\cdot(1,0,[1))\mathfrak{e}}_{n_{2}\hbox{\tiny{-times}}}=$
$=(s+p,s,[0))^{n_{2}}=$
$=(s+n_{2}p,s,[0)),$
@@
$(k(m+1),k(m+1),[0))$ $=(m+1,m+1,[0))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(s+pm,s+pm,[0))=$
$=(m,m,[1))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(m,m,[0))\mathfrak{e}=$
$=(km,km,[0)).$
@@
$(i,j,[0))\gamma_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[0))\gamma_{k_{1}}\delta_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[0))\delta_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[1))\delta_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}(i+1),k_{1}(j+1),[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
@@
$(i,j,[0))\delta_{k_{1}}\delta_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
@@
$(i,j,[1))\delta_{k_{1}}\delta_{k_{2}}$ $=(k_{1}(i+1),k_{1}(j+1),[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
@@
By $\mathfrak{e}_{\boldsymbol{0}}$ we denote the annihilating monoid endomorphism of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$, i.e., $(i,j,[p))\mathfrak{e}_{\boldsymbol{0}}=(0,0,[0))$ for all $i,j\in\omega$ and $p=0,1$. We put $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})=\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})\setminus\{\mathfrak{e}_{\boldsymbol{0}}\}$. Theorem 2 implies that $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ is a subsemigroup of $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
@@
    $\mathfrak{e}_{1}\mathscr{D}\mathfrak{e}_{2}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $\mathfrak{e}_{1}=\mathfrak{e}_{2}$ or there exists a positive integer $k$ such that $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\{\gamma_{k},\delta_{k}\}$;
@@
Since $\mathfrak{e}_{\boldsymbol{0}}$ is zero of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, the classes of equivalence of Green’s relations of non-zero elements of $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ in the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ coincide with their corresponding classes of equivalence in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, and moreover we have that
$\boldsymbol{L}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{R}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{H}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{D}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{J}_{\mathfrak{e}_{\boldsymbol{0}}}=\{\mathfrak{e}_{\boldsymbol{0}}\}$
in the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Acknowledgements
The authors acknowledge the referee for his/her comments and suggestions.